# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v3

@@ Abstract
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1], such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11]. The importance of time series forecasting cannot be overstated. There are also many methods for time series forecasting. The initial methods came from statistics, such as Autoregressive (AR) model, Moving Average (MA) model, Autoregressive Moving Average (ARMA) model, and Autoregressive integrated moving average (ARIMA) model [12]. Deep learning is a powerful method with wide applications in various disciplines [13, 14]. In recent years, time series prediction models based on neural networks have become a research hotspot. These neural networks have evolved from the simplest deep neural networks (DNN) [15]. Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on. Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [22, 23, 24].
There is uncertainty in time series data, and modeling uncertainty on time series is effective. Evidence theory is an uncertain reasoning framework based on mass function [25, 26, 27, 28]. Previous studies have shown that evidence theory can effectively utilize the uncertainty of time-series data and achieve positive application results [29, 30]. On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [31, 32]. Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [33]. Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion. TEFN is a lightweight neural network that considers the time and channel dimensions of multivariate time series as different sources of information in Figure 1.
@@ 1 Introduction
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20, 36]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. The contributions of this article are as follows:
@@
This paper is organized as follows. We begin by reviewing relevant literature in the Section 2. Next, in the Section 3, we delve into the architecture and underlying mathematical principles of TEFN. In the Section 4, we present a series of experiments that demonstrate TEFN’s superior performance. Finally, we conclude the paper by summarizing our key findings in the Section 5.
2 Related Work
2.1 Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [25, 26]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. All independent and mutually exclusive sample sets that exist in the system are called frames of discrete (FOD), represented as $\Omega$.
@@ 2.1 Dempster-Shafer Theory
2.2 Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making. The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (2) where $w$ and $b$ is weight and bias of the membership function. BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (3).
@@ 2.2 Basic Probability Assignment
$\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$ (3)
The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
2.3 Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [34, 37]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR [25, 26] mentioned in Section 2.1. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [38] in Equation (4), where $x$ is a target label and $e$ are combinations of all possible target labels.
$p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$ (4)
2.4 Linear Time Series Model
@@
TEFN shares similarities with Dlinear in that both models utilize linear regression for prediction. However, there are key differences:
  • DLinear employs a simpler decomposition method compared to TEFN, which may limit its ability to capture intricate temporal dynamics.
  • DLinear focuses on capturing linear relationships and trends in the data, while TEFN focuses on the mining of internal hidden relationships, so it strengthens predictions through dimensionality increase and uncertain reasoning.
  • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
3 Time Evidence Fusion Network
@@
3.1 Time Normalization and De-normalization
@@
3.2 Time Dimension Projection
@@
3.3 Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter.
This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation (10). Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation (10) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space. Figure 4 shows the visible process of BPA module.
@@ 3.3 Basic Probability Assignment Module
By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions. This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications.
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation (10), we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including:
  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation (3), capturing the likelihood of each data point belonging to different membership levels. The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation (11). The sigmoid function can capture both gradual and abrupt changes in membership values.
    $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$ (11)
  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation (12) where $g(x)$ is a non-linear function. This approach is useful for capturing complex patterns and relationships within the data.
    $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$ (12)
By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data. This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
3.4 Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation (13). Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
@@ 3.4 Expectation Fusion
The fusion operation in Equation (13) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
@@
4 Experiments
@@
4.1 Dataset
Our research employs the TEFN for the challenging task of long-term time series prediction. To ensure the model’s effectiveness and robustness across diverse scenarios, we meticulously selected five representative and extensively used datasets. These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance. The selected datasets are:
  • Electricity [44]: This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time. It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
  • ETT (4 subsets) [19]: This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges. It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
  • Exchange [45]: This dataset captures exchange rate fluctuations between major currencies. It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
  • Traffic [46]: This dataset contains traffic flow data collected from sensors in a city. It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
  • Weather [47]: This dataset includes historical weather data, encompassing various meteorological parameters. It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table I.
@@ 4.1 Dataset
4.2 Comparative Baseline
To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models. These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities. The selected baseline models are as follows:
  • Transformer-based Models:
    • iTransformer (ICLR 2024) [48]: A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
    • PatchTST (ICLR 2023) [36]: This model divides the time series into patches and applies transformer-based operations on each patch independently. It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
    • Crossformer (ICLR 2023) [20]: This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
    • FEDformer (ICML 2022) [52]: A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
    • ETSformer [51]: A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
    • Stationary (NeurIPS 2022) [42]: A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
    • Autoformer (NeurIPS 2021) [45]: A self-supervised transformer model that learns to forecast future values without relying on explicit labels. It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
  • Linear Models:
    • Rlinear (ICLR 2024) [49]: This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data. It is effective at capturing periodic patterns and works well with longer input sequences.
    • Dlinear (AAAI 2023) [39]: A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data. It serves as a simple yet effective baseline for comparison.
    • LightTS [40]: A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
  • Other Models:
    • TiDE (ICLR 2023) [50]: A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
    • TimesNet (ICLR 2023) [43]: A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.

### v4

@@ Abstract
I Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1] such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11]. The importance of time series forecasting cannot be overstated. There are many methods for time series forecasting. The initial methods came from statistics, such as Autoregressive integrated moving average (ARIMA) model [12]. Deep learning is a powerful method with wide applications in various disciplines [13, 14]. In recent years, time series prediction models based on neural networks have become a research hotspot. These neural networks have evolved from the simplest deep neural networks (DNN) [15]. Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on [22, 23]. Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [24, 25, 26].
There is uncertainty in time series data, and modeling uncertainty on time series is effective. Evidence theory is an uncertain reasoning framework based on mass function [27, 28, 29, 30]. Previous studies have shown that evidence theory can effectively utilize the uncertainty of time series data and achieve positive application results [31, 32]. On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [33, 34]. Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [35]. Furthermore, the essence of evidence theory lies in using mass functions to model variables. A mass function, as a generalized form of probability, can directly represent the uncertainty among different arguments that is challenging to achieve with traditional probability models. Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion. TEFN is a lightweight neural network that considers the time and channel dimensions of multivariate time series as different sources of information in Figure 1.
@@ I Introduction
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20, 38]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. TEFN holds significant value and potential impact in various practical application scenarios. In the energy sector, such as power systems, accurate electricity consumption forecasting is crucial for the reliable operation, management, and planning of the power grid [39, 40]. With large scale power consumption data, traditional forecasting methods may face high computational complexity and long training times. TEFN, however, can efficiently process millions of time series, enabling power companies to optimize power generation and distribution, reducing energy waste and costs. In the context of AIOps, resource forecasting and optimization require accurate predictions to ensure the stable operation of cloud systems [41]. As cloud platforms manage millions of time series data related to resource usage, low complexity and high efficiency forecasting methods are essential [42]. TEFN can provide precise resource usage forecasts, helping cloud operators to allocate resources effectively, reduce operational costs, and improve service reliability.
The contributions of this article are as follows:
@@
This paper is organized as follows. We begin by reviewing relevant literature in the Section II. Next, in the Section III, we delve into the architecture and underlying mathematical principles of TEFN. In the Section IV, we present a series of experiments that demonstrate TEFN’s superior performance. Finally, we conclude the paper by summarizing our key findings in the Section V.
II Related Work
II-A Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [27, 28]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. All independent and mutually exclusive sample sets that exist in the system are called frames of discernment (FOD), represented as $\Omega$.
@@ II-A Dempster-Shafer Theory
Here is an example of an intuitive explanation of the mass function. Suppose there are two unidentifiable features $A$ and $B$ in the time series. There are different ways $1$ and $2$ to identify whether it is $A$ or $B$ (For example, whether the increase in time series values is due to a trend or a cycle.). Therefore, we obtain the results in in Equation (2), and both ways are able to distinguish the feature of $A$, but there is also a certain degree that couldn’t distinguish between $A$ and $B$. It should be noted that if modeling is done using probabilities, the inability to distinguish between $A,B$ cannot be directly expressed. The usual approach is to declare a new argument $C$ as indistinguishable between $A$ and $B$. But at this time, $C$ will be independent of $A$ and $B$, and the connection between propositions will be lost. DSR can unify the opinions of the two methods $m_{1},m_{2}$ and obtain a comprehensive opinion $m$ in Equation (2).
[…]
II-B Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making.
The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (3) where $w$ and $b$ is weight and bias of the membership function. This kind of membership is called the triangular membership function, assuming that the membership is linearly related to the random variable.
@@ II-B Basic Probability Assignment
An intuitive example is shown in Figure 3. There are two fuzzy sets, which are assumed to represent the categories of random variables. We randomly take a sample $x_{i}$, and at this time, the membership function reflects the degree to which $x_{i}$ belongs to this fuzzy set, which can be understood as the degree of belonging to this category. This is also different from probability. The membership degree is bounded and can only reflect the extent to which it belongs to the current fuzzy set. Therefore, when multiple fuzzy sets appear, they need to be normalized in order to be converted into a probability distribution or a mass function. And the mass function can increase the description of uncertainty. In Figure 3, since the degree of belonging to the first category is greater, we believe that the degree of belonging to the second category actually reflects that the two fuzzy sets are not clearly distinguishable.
Figure 3: The conversion between fuzzy membership and mass function.
BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (taking probability density function as a membership function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (4).
[…]
The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information. On the one hand, the existence of fuzzy sets allows elements to partially belong to a set, which allows for a more continuous attribution, rather than the set only belonging to this set or not belonging to this set.
II-C Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [36, 43]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR [27, 28] mentioned in Section II-A. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [44] in Equation (5), where $x$ is a target label and $e$ are combinations of all possible target labels.
$p(x)=\sum_{x\in e}\frac{m(e)}{|e|}$ (5)
II-D Linear Time Series Model
@@
III Time Evidence Fusion Network
@@
III-A Time Normalization and De-normalization
@@
III-B Time Dimension Projection
@@
III-C Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter. It can be considered that the BPA operation can be regarded as an increase in dimension, and the size of the new dimension is the same as the power set. This operation expands the different representations of the original elements on the power set. The simplest form of the generating function is a linear function, which can be understood as a fuzzy membership function. Of course, it can also be a nonlinear function. In neural networks, a nonlinear layer is usually added after the linear transformation.
This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation (11). Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation (11) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function. The intuitive understanding of Equation (11) is that the different values $j$ on the input data dimension $D$ are expanded into power sets, where $k$ corresponds to different elements in the power set. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
@@ III-C Basic Probability Assignment Module
Figure 5 shows the visible process of BPA module in time dimension. It should be noted that this is not matrix multiplication, but a multiplication of the same number of different parameters for each element, and the results are combined into a new dimension. The values in the Figure 5 corresponding to different time steps are extended to the mass of the fuzzy set memship function measure in the event space. These membership functions are shared among different channels. Therefore, the size of the newly added dimension is $2^{|S|}$, reflecting the mass function of different events. The operation on the channel dimension is symmetrical, and it shares parameters for different time steps.
By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions. This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications. Unlike the Bayesian layer and the probability layer, BPA considers imprecise probabilities, allowing for the uncertainty of the probability modeling itself, rather than using probabilities as a basis to estimate uncertainty. An intuitive example is the DSR case in Section II-A.
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation (11), we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including ReLU Function, Tanh Function, Piecewise Functions, etc. By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data. This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
III-D Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation (12). Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension. The fusion operation in Equation (12) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
@@ III-D Expectation Fusion
III-E Comparison With Other Lightweight Forecasting Models
Common lightweight models, such as DLinear [45] and LightTS [46], focus on the characteristics of the time series itself through different approaches. The main differences between TEFN and other lightweight models are as follows:
  • Theoretical Foundation: While DLinear employs seasonal-trend decomposition and LightTS uses sampling techniques, evidence-based framework of TEFN fundamentally differs by modeling uncertainty through mass functions rather than direct feature extraction. This allows TEFN to handle ambiguous temporal patterns that challenge traditional decomposition methods.
  • Feature Engineering: DLinear requires explicit decomposition of trends and seasonality, making it sensitive to decomposition accuracy. TEFN eliminates this need through its BPA module’s automatic uncertainty quantification, as shown in Equation (12).
  • Computational Efficiency: Though all models emphasize efficiency, TEFN achieves parameter reduction through expectation fusion rather than architectural simplification. Expectation fusion directly adds the results of time BPA and channel BPA, without multiplication operation that the acceleration mode is similar to [50], and without introducing new modules. The parameter amount occupied in fusion operation is 0. Also, this will become a potential limitation, because in the engineering application scenarios of TEFN, due to the different characteristics of the data, other fusion methods may be used, such as concat and then use the linear link layer for fusion, or even use the attention mechanism with larger parameters for fusion.
IV Experiments
@@
IV-A Dataset
Our research employs the TEFN for the challenging task of long-term time series prediction. To ensure the model’s effectiveness and robustness across diverse scenarios, we meticulously selected five representative and extensively used datasets. These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance. The datasets are Electricity [51], ETT (4 subsets) [19], Exchange [52], Traffic [53], and Weather [54]. Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table I.
@@ IV-A Dataset
IV-B Comparative Baseline
To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models. These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities. The selected baseline models are as follows: Transformer-based Models, including iTransformer (ICLR 2024) [55], Crossformer (ICLR 2023) [20], FEDformer (ICML 2022) [59], ETSformer [58], Stationary (NeurIPS 2022) [48], and Autoformer (NeurIPS 2021) [52]; Linear Models, such as Rlinear (ICLR 2024) [56], DLinear (AAAI 2023) [45], and LightTS [46]; and Other Models, like TiDE (ICLR 2023) [57] and TimesNet (ICLR 2023).